OpenStudy (hockeychick23):

What is _10C_8 ? (The 10 and 8 are subscripts next to the C, thanks!)

geerky42 (geerky42):

\(\displaystyle_{10}C_{8} = \binom{10}{8} = \dfrac{10!}{8!~2!} = \dfrac{10\cdot9\cdot8\cdot7\cdot6\cdot5\cdot4\cdot3\cdot2}{8\cdot7\cdot6\cdot5\cdot4\cdot3\cdot2\cdot2} \\~\\~\\\quad\quad~= \dfrac{10\cdot9\cdot\cancel{8}\cdot\cancel{7}\cdot\cancel{6}\cdot\cancel{5}\cdot\cancel{4}\cdot\cancel{3}\cdot\cancel{2}}{\cancel{8}\cdot\cancel{7}\cdot\cancel{6}\cdot\cancel{5}\cdot\cancel{4}\cdot\cancel{3}\cdot\cancel{2}\cdot2} \\~\\~\\\quad\quad~= \dfrac{10\cdot9}{2} \\~\\~\\\quad\quad~= \cdots~?\)

OpenStudy (hockeychick23):

90/2

OpenStudy (hockeychick23):

(45)

geerky42 (geerky42):

Yeah, what is \(\dfrac{90}{2}\) equal to?

geerky42 (geerky42):

yep answer is 45

OpenStudy (hockeychick23):

thanks!

geerky42 (geerky42):

Do you understand what I did?

OpenStudy (hockeychick23):

Yea i do now, i looked at the work you did at the top with the descending numbers, thanks!

geerky42 (geerky42):

Ok glad I helped!